OpenStudy (anonymous):

Which of the following units is the SI unit of velocity? a. meter c. meter per second b. meter·second d. second per meter

OpenStudy (konradzuse):

Lets look at it this way... What is the formula for Velocity?

OpenStudy (konradzuse):

???

OpenStudy (anonymous):

meters/second right?

OpenStudy (konradzuse):

yes, do you know the formula? Velocity =????

OpenStudy (anonymous):

isn't it just m/s^2

OpenStudy (konradzuse):

nope, that is acceleration.

OpenStudy (anonymous):

oh then no i don't know it.

OpenStudy (konradzuse):

The Formula is Velocity = Distance/Time.

OpenStudy (anonymous):

oh, thank you!

OpenStudy (konradzuse):

Formulas are very important to remember in Physics... What is Distance Measured in? Time?

OpenStudy (konradzuse):

Acceleration is the change in Velocity/ change in Time.

OpenStudy (anonymous):

meters and seconds usually. I just have this e2020 class and my teacher is just making me take the tests and not giving me the formulas in the direct instruction

OpenStudy (konradzuse):

yes, that is correct... Hmm bad teachers are no good :(. You can try complaning and maybe getting a class switch? A = V/T = m/s/s = m/s^2

OpenStudy (anonymous):

Thank You.

OpenStudy (konradzuse):

Yup, you can tag me if you have anymore questions.. Medal appreciated :).